OpenStudy (anonymous):

The maximum value of f(x) = -e^x + x is:

OpenStudy (anonymous):

A) 0 B) 2 C) ln 2 D) -1 E) 1

OpenStudy (anonymous):

Diff it and put equal to 0

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

f'(x) = -e^x

OpenStudy (anonymous):

What about the other bit?

OpenStudy (anonymous):

+ 1

OpenStudy (anonymous):

f'(x) = -e^x + 1

OpenStudy (anonymous):

So put it equal to O

OpenStudy (anonymous):

0 = -e^x + 1

OpenStudy (anonymous):

x=?

OpenStudy (anonymous):

the e is euler's constant right?

OpenStudy (anonymous):

Euler's constant is something else, sometimes it's called Euler's number

OpenStudy (anonymous):

Anyway, u got your answer, right?

OpenStudy (anonymous):

nope. can you take it from here?

OpenStudy (anonymous):

e is the number 2.7... etc. Subtract 1 from the above equation, you will get: -1 = -e^x => e^x = 1 Take the natural log on both sides: x = ln(1) = 0 :-)

OpenStudy (anonymous):

So now sub that in your original to get the max value.

OpenStudy (anonymous):

-1 yay! thanks you guys.